A prisoner is considering his chances of escape. From his cell, there are 3 possible exits he could choose and if he chooses a particular exit, the probability he will escape is 0.4; 0.3 and 0.4 respectively. If the prisoner picks an exit at random, what is the probability that the prisoner will make a successful escape (to 3 decimal places)?


1
Expert's answer
2022-05-03T14:35:32-0400
"P(1)P(Yes|1)+P(2)P(Yes|2)+P(3)P(Yes|3)"

"=\dfrac{1}{3}(0.4)+\dfrac{1}{3}(0.3)+\dfrac{1}{3}(0.4)=\dfrac{11}{30}\approx0.367"